Chapter:The Elimination Method
2m=2-9n
21n=4-6m

Answer by checkley75(3666)

-9N=-2-2M NOW MULTIPLY THIS EQUATION BY -3 & ADD.
21N=4-6M
27N=6+6M
----------------------------------
48N=10
N=10/48
N=5/24 ANSWER. NOW SUBSTITUTE 5/24 FOR N & SOLVE FOR M
2M=2-9*5/24
2M=2-45/24
2M=(48-45)/24
2M=3/24
2M=1/8
M=1/8*1/2
M=1/16 ANSWER.